OpenStudy (anonymous):

Kaihi can eat 32 lollies in one hour. Her sister can eat the same amount in 3 hours Howlong will it take both girls to eat 32 lollies?

OpenStudy (anonymous):

45minutes

OpenStudy (anonymous):

= 3/4 hr

OpenStudy (anonymous):

@AbhimanyuPudi is right 45 min

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Kaihi eats 32 lollies in 1hr..her sis eats 32 in 3hrs which means her sis eats 32/3 in 1hr... So, in 1hr Kaihi eats 32 while her sis eats 32/3 If kaihi eats x lollies in a partcular time, then her sis eats x/3 lollies in the same time... Now, according to question x + x/3 = 32 By solving you get, x=24 So, Kaihi eats 24 lollies and her sister eats 8. Since we need time in which they eat this way: Kaihi eats 32 in 1hr So, she can eat 24 in 24/32 hr = 3/4 hr = 45min

OpenStudy (anonymous):

welcome:)